OpenStudy (anonymous):

Suppose you drive to a picnic spot at 20 mph,then how fast you must go while returning back home on the same route such that your average speed becomes 40 mph ? Brain Teaser

OpenStudy (dtan5457):

oh you and all your brain teasers :D

Nnesha (nnesha):

40 is total so

OpenStudy (anonymous):

Brain teasers are really fun

Nnesha (nnesha):

start with 20 :O_O O_O ^_^

Nnesha (nnesha):

then what ?? :P:D

Nnesha (nnesha):

can i ask one brain teaser question ??

OpenStudy (anonymous):

Sure go ahead

Nnesha (nnesha):

1+1 = 11 how

OpenStudy (dtan5457):

60 mph? 60+20=80 80/2=40? lol it looks simple when i do it like that

Nnesha (nnesha):

if you don't know then ask me HOW ? :) ;)

OpenStudy (dtan5457):

i dont have a clue

OpenStudy (dtan5457):

maybe the guy thought adding numbers was just to combine

OpenStudy (dtan5457):

and put them together

Nnesha (nnesha):

@tbing5 wbu shamart :)